2 Al + 3 CuCl₂ → 2 AICI3 + 3 Cu



Molar masses: 27.0 g/mol Al,


134.5 g/mol CuCl2, 63.5 g/mol Cu



A: How many grams of CuCl₂ are


necessary to consume 10 g Al?


B: How many grams of Cu will be


produced?

1
Expert's answer
2022-06-03T05:00:05-0400

Solution:

The molar mass of Al is 27.0 g/mol

Therefore,

Moles of Al = (10 g Al) × (1 mol Al / 27.0 g Al) = 0.3704 mol Al


Balanced chemical equation:

2Al + 3CuCl2 → 2AICI3 + 3Cu

According to stoichiometry:

2 mol of Al react with 3 mol of CuCl2

Thus, 0.3704 mol of Al react with:

(0.3704 mol Al) × (3 mol CuCl2 / 2 mol Al) = 0.5556 mol CuCl2


The molar mass of CuCl2 is 134.5 g/mol

Therefore,

Mass of CuCl2 = (0.5556 mol CuCl2) × (134.5 g CuCl2 / 1 mol CuCl2) = 74.7282 g CuCl2 = 74.73 g CuCl2

Thus, 74.73 grams of CuCl2 are necessary to consume 10 g Al


According to stoichiometry:

2 mol of Al produce 3 mol of Cu

Thus, 0.370 mol of Al produce:

(0.3704 mol Al) × (3 mol Cu / 2 mol Al) = 0.5556 mol Cu


The molar mass of Cu is 63.5 g/mol

Therefore,

Mass of Cu = (0.5556 mol Cu) × (63.5 g Cu / 1 mol Cu) = 35.2806 g Cu = 35.28 g Cu

Thus, 35.28 grams of Cu will be produced


Answer:

A: 74.73 grams of CuCl2

B: 35.28 grams of Cu